OpenStudy (anonymous):

Statistics problem: how would I find the probability of at least 3 in 5 people being left handed? Assuming for this problem that the probability of being right handed is .90

OpenStudy (anonymous):

At least three means three, four, or five. Calculate the probability that three people are left handed, then the probability that four people are left handed, then the probability that five people are left handed, and add them all together.

OpenStudy (dumbcow):

to calculate those probabilities use the binomial theorem \[P(X=k) = \left(\begin{matrix}n \\ k\end{matrix}\right)p^{k}(1-p)^{n-k}\] here n=5 , p = 0.1